4V . . . . 8V.. ........ .+. . . . VCE + VBET ... .. For the given above circuit. VBE = 0.7V. If a = 0.986 find ß of the BJT..

Using the above parameters, a = 0.986, RC = 212, RE=2.0 k 2, design RB such that VE = 2V.

Using the above parameters,Find VCE.

Find B, RB, VCE
Answer #1

Be a la 70.4286 Vezav. RE OKIRC=2K 26 E = 2 = mA kre for loop 4 - RB IB IVBE +2 IB = IE=0.014mA RB = 4.0.7-2 0.014m = 92.8571